OpenStudy (anonymous):

For 𝑓(𝑥)=17𝑥+3, find f(2).

OpenStudy (anonymous):

please help me with as many as you can! i dont want to be suspended :(

OpenStudy (gabylovesyou):

so just substitute every x for 2 :) f(x) = 17x+3 turns into f(2) = 17 * 2 + 3 in PEMDAS you should multiply first so 17 * 2 = 34 f(2) = 34 + 3 34 + 3 = 37 so f(2) = 37 :)